# Diederich–Fornæss index and global regularity of the complex Green operator: domains with comparable Levi eigenvalues

### v1

@@ 4. The main estimate
(4.49) $\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\left|\alpha_{\eta}(L_{\mathcal{P}^{+,\eta}u}^{K})\right|^{2}\\ \lesssim(1-\eta)\Big(\left\|\bar{\partial}_{b}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}(\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\left\|\bar{\partial}_{b}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}+\|\overline{\partial}_{b}u\|^{2}+\|\overline{\partial}_{b}^{*}u\|^{2}\Big)\\ +C_{\eta}\|u\|_{-1}^{2}\;,$
@@
note that $(1-\eta)^{2}\leq(1-\eta)$. The various terms on the right hand side of (4.49) are estimated as follows. First,
@@
We have again used maximal estimates and $\left\|u\right\|^{2}\lesssim\left\|\bar{\partial}_{b}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}u\right\|^{2}$. In the terms in the first line of the right hand side of (4.49) we commute $\mathcal{P}^{+,\eta}$ with the respective differential operators. Because $(\mathcal{P}^{+,\eta}-\mathcal{P}^{+})$ is infinitely smoothing ($\chi^{+,\eta}$ and $\chi^{+}$ differ by a compactly supported symbol), and $\overline{\partial}_{b},\mathcal{P}^{+}]$ is of order zero (and does not depend on $\eta$), we can prove estimates that are uniform in $\eta$ (modulo a weak term):
(4.51) $\overline{\partial}_{b}\mathcal{P}^{+,\eta}u\|^{2}\lesssim\|\overline{\partial}_{b}\mathcal{P}^{+}u\|^{2}+\|\overline{\partial}_{b}(\mathcal{P}^{+,\eta}-\mathcal{P}^{+})u\|^{2}\\ \lesssim\|\mathcal{P}^{+}\overline{\partial}_{b}u\|^{2}+\|[\overline{\partial}_{b},\mathcal{P}^{+}]u\|^{2}+\|(\mathcal{P}^{+,\eta}-\mathcal{P}^{+})u\|_{1}^{2}\\ \lesssim\|\overline{\partial}_{b}u\|^{2}+\|u\|^{2}+C_{\eta}\|u\|_{-1}^{2}\lesssim\|\overline{\partial}_{b}u\|^{2}+\|\overline{\partial}_{b}^{*}u\|^{2}+C_{\eta}\|u\|_{-1}^{2};.$
@@
(4.52) $\begin{split}\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\left|\alpha_{\eta}(L_{\mathcal{P}^{+,\eta}u}^{K})\right|^{2}&\lesssim(1-\eta)\Big(\left\|\bar{\partial}_{b}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}u\right\|^{2}\Big)+C_{\eta}\left\|u\right\|_{-1}.\end{split}$
@@ 5. Proof of Theorem 1
We first prove $(1)$ for $\min\{q_{0},n-1-q_{0}\}\leq q\leq\max\{q_{0},n-1-q_{0}\}$. To simplify notation from section 3, now that there is an additional subscript $\eta$, we replace the double subscripts $\rho_{\eta}$ by the simple subscript $\eta$. That is $T_{\rho_{\eta}}$ becomes $T_{\eta}$, $\alpha_{\rho_{\eta}}$ becomes $\alpha_{\eta}$, and so on. Note that these fields and forms live in a neighborhood of $b\Omega=b\Omega_{\eta}$ and so can be used on $b\Omega_{\eta,\delta}$. Constants used in the proof may vary form one occurrence to the next, but are independent of $\eta$, unless they come with a subscript $\eta$. (Since $k$ is fixed throughout the argument, we do not indicate the dependence on $k$.) The estimates below should be understood as ‘there is $\delta_{\eta}$ so that the estimate holds uniformly in $\delta$ for $\delta<\delta_{\eta}$’. In order not to clutter the notation, we do not use subscripts $(\eta,\delta)$ on the norms.
@@
Recall from section 3 the vector fields $L_{u}^{J}$, $|J|=q$, associated to a (0,q)(0,q)–form, and the fact that although these vector fields are not invariant under a change form the global Euclidean frame to a special boundary frame, the quantity on the left hand side of (4.1) in Proposition 1 is, up to a constant factor. Further recall from (4.42)–(4.43) that if $J^{\prime}\subset J$ is a (q1)(q-1)–tuple with $j\notin J^{\prime}$, then $L_{u_{J}\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}}}^{J^{\prime}}=u_{J}\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}}$. Finally, $|\alpha_{\eta}(\overline{L_{j}})|=|\alpha_{\eta}(L_{j})|$, because $\alpha_{\eta}$ is real. With these facts in mind, the commutator formula (3.3) gives
(5.12) $\big\|\chi_{s}[\overline{\partial}_{b},T_{\eta}^{k}]u\big\|^{2}\lesssim\sum_{j\notin J}\sideset{}{{}^{\prime}}{\sum}_{J}\|\alpha_{\eta}(\chi_{s}T_{\eta}^{k}u_{J}L_{j})\|^{2}+C_{\eta}\big(\|\overline{\partial}_{b}u\|_{k-1}^{2}+\|\overline{\partial}_{b}^{*}u\|_{k-1}^{2}+\|u\|_{k-1}^{2}\big)\\ =\sum_{j\notin J}\sum_{J}\int_{b\Omega_{\eta,\delta}}\big|\alpha_{\eta}(L^{J^{\prime}}_{\chi_{s}T_{\eta}^{k}u_{J}(\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})})\big|^{2}+C_{\eta}\big(\|\overline{\partial}_{b}u\|_{k-1}^{2}+\|\overline{\partial}_{b}^{*}u\|_{k-1}^{2}+\|u\|_{k-1}^{2}\big)\;.$
@@
(5.13) $\sum_{j\notin J}\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega_{\eta,\delta}}\big|\alpha_{\eta}(L^{J^{\prime}}_{\chi_{s}T_{\eta}^{k}u_{J}(\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})})\big|^{2}\leq\sum_{j\notin J}\sideset{}{{}^{\prime}}{\sum}_{J}\sideset{}{{}^{\prime}}{\sum}_{|K|=q-1}\int_{b\Omega_{\eta,\delta}}\big|\alpha_{\eta}(L^{K}_{\chi_{s}T_{\eta}^{k}u_{J}(\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})})\big|^{2}\\ \lesssim(1-\eta)\sum_{j\notin J}\sideset{}{{}^{\prime}}{\sum}_{J}\bigg(\|\overline{\partial}_{b}(\chi_{s}T_{\eta}^{k}u_{J}(\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}}))\|^{2}+\|\overline{\partial}_{b}^{*}(\chi_{s}T_{\eta}^{k}u_{J}(\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}}))\|^{2}\bigg)\\ +C_{\eta}\sum_{j\notin J}\sideset{}{{}^{\prime}}{\sum}_{J}\|\chi_{s}T_{\eta}^{k}u(\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})\|_{-1}^{2}$
@@
(5.14) $\sum_{j\notin J}\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega_{\eta,\delta}}\big|\alpha_{\eta}(L^{J^{\prime}}_{\chi_{s}T_{\eta}^{k}u_{J}(\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})})\big|^{2}\\ \lesssim(1-\eta)\big(\|\overline{\partial}_{b}(\chi_{s}T_{\eta}^{k}u)\|^{2}+\|\overline{\partial}_{b}^{*}(\chi_{s}T_{\eta}^{k}u)\|^{2}+\|\chi_{s}T_{\eta}^{k}u\|^{2}\big)+C_{\eta}\|u\|_{k-1}^{2}\\ \lesssim(1-\eta)\big(\|\chi_{s}\overline{\partial}_{b}(T_{\eta}^{k}u)\|^{2}+\|\chi_{s}\overline{\partial}_{b}^{*}(T_{\eta}^{k}u)\|^{2}+\|\hat{\chi}_{s}T_{\eta}^{k}u\|^{2}\big)+C_{\eta}\|u\|_{k-1}^{2}\;,$
@@
Next, we similarly address the commutator term with $\overline{\partial}_{b}^{*}$ on the right hand side of (5.11). The argument is actually simpler; because $\sum_{j}\chi_{s}(T_{\eta}^{k}u_{jS})L_{j}=\sum_{j}(\chi_{s}T_{\eta}^{k}u)_{jS}L_{j}=L^{s}_{\chi_{s}T_{\eta}^{k}u}$, there is no need for ‘auxiliary’ forms. Formula (3.4) and Proposition 1 give
(5.16) $\|\chi_{s}[\overline{\partial}_{b}^{*},T_{\eta}^{k}]u\|^{2}\\ \lesssim\sideset{}{{}^{\prime}}{\sum}_{S}\big\|\sum_{j}\chi_{s}\alpha_{\eta}(L_{j})T_{\eta}^{k}u_{jS}\big\|^{2}+C_{\eta}\big(\|\overline{\partial}_{b}u\|_{k-1}^{2}+\|\overline{\partial}_{b}^{*}u\|_{k-1}^{2}+\|u\|_{k-1}^{2}\big)\\ =\sideset{}{{}^{\prime}}{\sum}_{S}\|\alpha_{\eta}(L^{S}_{\chi_{s}T_{\eta}^{k}u})\|^{2}+C_{\eta}\big(\|\overline{\partial}_{b}u\|_{k-1}^{2}+\|\overline{\partial}_{b}^{*}u\|_{k-1}^{2}+\|u\|_{k-1}^{2}\big)\\ \lesssim(1-\eta)\big(\|\overline{\partial}_{b}(\chi_{s}T_{\eta}^{k}u)\|^{2}+\|\overline{\partial}_{b}^{*}(\chi_{s}T_{\eta}^{k}u)\|^{2}\big)\;\;\;\;\;\;\;\;\;\;\;\;\;\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+C_{\eta}\big(\|\overline{\partial}_{b}u\|_{k-1}^{2}+\|\overline{\partial}_{b}^{*}u\|_{k-1}^{2}+\|u\|_{k-1}^{2}\big)\\ \;\;\;\;\;\;\;\;\;\lesssim(1-\eta)\big(\|\chi_{s}\overline{\partial}_{b}(T_{\eta}^{k}u)\|^{2}+\|\chi_{s}\overline{\partial}_{b}^{*}(T_{\eta}^{k}u)\|^{2}+\|\hat{\chi}_{s}T_{\eta}^{k}u\|^{2}\big)\\ +C_{\eta}\big(\|\overline{\partial}_{b}u\|_{k-1}^{2}+\|\overline{\partial}_{b}^{*}u\|_{k-1}^{2}+\|u\|_{k-1}^{2}\big)\;.$

### v2

@@ 4. The main estimate
(4.46) $\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{J}\left\|(\overline{L}_{j}h_{\eta})\mathcal{P}^{+,\eta}u_{J}\right\|^{2}\\ \lesssim(1-\eta)\Big(\left\|\bar{\partial}_{b}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u\right\|^{2}+\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{J}\Big(\left\|\bar{\partial}_{b}^{\star}(\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\\ \;\;\;\;\;\;\;\;\;+\left\|\bar{\partial}_{b}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\Big)\Big)+\|u\|^{2}+C_{\eta}\|u\|_{-1}^{2}\;.$
We have used that $\|\mathcal{P}^{+,\eta}u\|^{2}\lesssim\|u\|^{2}$ and $\left\|u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}}\right\|_{-1}^{2}\lesssim\|u\|_{-1}^{2}$. Starting with the first inequality in (4.41) and using (4.46) gives us
(4.47) $\left\|\bar{\partial}_{b}h_{\eta}\wedge\mathcal{P}^{+,\eta}u\right\|^{2}\\ \lesssim(1-\eta)\Big(\left\|\bar{\partial}_{b}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u\right\|^{2}+\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{J}\big(\left\|\bar{\partial}_{b}^{\star}(\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\\ \;\;\;\;\;\;\;\;\;+\left\|\bar{\partial}_{b}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\big)\Big)+\|u\|^{2}+C_{\eta}\|u\|_{-1}^{2}\;.$
Combining (4.39), (4.40), and (4.47), and using $(1-\eta)^{2}\leq(1-\eta)$ and (2.2), we get
(4.48) $\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\left|\alpha_{\eta}(\tilde{L}_{\mathcal{P}^{+,\eta}u}^{K})\right|^{2}\\ \lesssim(1-\eta)\Big(\left\|\bar{\partial}_{b}\mathcal{P}^{+,\eta}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}\mathcal{P}^{+,\eta}u\right\|^{2}+\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{J}\Big(\left\|\bar{\partial}_{b}^{\star}(\mathcal{P}^{+,\eta}u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\left\|\bar{\partial}_{b}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}(u_{J}\bar{\omega}_{j}\wedge\bar{\omega}_{J^{\prime}})\right\|^{2}\Big)+\|\overline{\partial}_{b}u\|^{2}+\|\overline{\partial}_{b}^{*}u\|^{2}\Big)\\ +C_{\eta}\|u\|_{-1}^{2}\;.$
@@
The various terms on the right hand side of (4.48) are estimated as follows. First,
@@
We have again used maximal estimates and (2.2). In the terms in the first line of the right hand side of (4.48) we commute $\mathcal{P}^{+,\eta}$ with the respective differential operators. Because $(\mathcal{P}^{+,\eta}-\mathcal{P}^{+})$ is infinitely smoothing ($\chi^{+,\eta}$ and $\chi^{+}$ differ by a compactly supported symbol), and $\overline{\partial}_{b},\mathcal{P}^{+}]$ is of order zero (and does not depend on $\eta$), we can prove estimates that are uniform in $\eta$ (modulo a weak term):
(4.50) $\overline{\partial}_{b}\mathcal{P}^{+,\eta}u\|^{2}\lesssim\|\overline{\partial}_{b}\mathcal{P}^{+}u\|^{2}+\|\overline{\partial}_{b}(\mathcal{P}^{+,\eta}-\mathcal{P}^{+})u\|^{2}\\ \lesssim\|\mathcal{P}^{+}\overline{\partial}_{b}u\|^{2}+\|[\overline{\partial}_{b},\mathcal{P}^{+}]u\|^{2}+\|(\mathcal{P}^{+,\eta}-\mathcal{P}^{+})u\|_{1}^{2}\\ \lesssim\|\overline{\partial}_{b}u\|^{2}+\|u\|^{2}+C_{\eta}\|u\|_{-1}^{2}\lesssim\|\overline{\partial}_{b}u\|^{2}+\|\overline{\partial}_{b}^{*}u\|^{2}+C_{\eta}\|u\|_{-1}^{2}.$
@@
(4.51) $\begin{split}\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\left|\alpha_{\eta}(\tilde{L}_{\mathcal{P}^{+,\eta}u}^{K})\right|^{2}&\lesssim(1-\eta)\Big(\left\|\bar{\partial}_{b}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}u\right\|^{2}\Big)+C_{\eta}\left\|u\right\|_{-1}.\end{split}$
@@ 5. Proof of Theorem 1
We first prove $(1)$ for $\min\{q_{0},n-1-q_{0}\}\leq q\leq\max\{q_{0},n-1-q_{0}\}$. Note that $T_{\eta}$ and $\alpha_{\eta}$ live in a neighborhood of $b\Omega=b\Omega_{\eta}$ and on $b\Omega_{\eta,\delta}$ agree with $T_{(\rho_{\eta}+\delta)}$ and $\alpha_{(\rho_{\eta}+\delta)}$, respectively. Constants used in the proof may vary form one occurrence to the next, but are independent of $\eta$, unless they come with a subscript $\eta$. (Since $k$ is fixed throughout the argument, we do not indicate the dependence on $k$.) The estimates below should be understood as ‘there is $\delta_{\eta}$ so that the estimate holds uniformly in $\delta$ for $\delta<\delta_{\eta}$’. In order not to clutter the notation, we do not use subscripts $(\eta,\delta)$ on the norms.
@@
Recall from section 3 the vector fields $L_{u}^{J}$, $|J|=q$, associated to a (0,q)(0,q)–form, and the fact that although these vector fields are not invariant under a change form the global Euclidean frame to a special boundary frame, the quantity on the left hand side of (4.1) in Proposition 1 is, up to a constant factor. Further recall from (4.41)–(4.42) that if $J^{\prime}\subset J$ is a (q1)(q-1)–tuple with $j\notin J^{\prime}$, then $\tilde{L}_{u_{J}\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}}}^{J^{\prime}}=u_{J}\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}}$. Finally, $|\alpha_{\eta}(\overline{L_{j}})|=|\alpha_{\eta}(L_{j})|$, because $\alpha_{\eta}$ is real. With these facts in mind, the commutator formula (3.3) gives
(5.12) $\big\|\chi_{s}[\overline{\partial}_{b},T_{\eta}^{k}]u\big\|^{2}\lesssim\sum_{j\notin J}\sideset{}{{}^{\prime}}{\sum}_{J}\|\alpha_{\eta}(\chi_{s}T_{\eta}^{k}u_{J}L_{j})\|^{2}+C_{\eta}\big(\|\overline{\partial}_{b}u\|_{k-1}^{2}+\|\overline{\partial}_{b}^{*}u\|_{k-1}^{2}+\|u\|_{k-1}^{2}\big)\\ =\sum_{j\notin J}\sum_{J}\int_{b\Omega_{\eta,\delta}}\big|\alpha_{\eta}(\tilde{L}^{J^{\prime}}_{\chi_{s}T_{\eta}^{k}u_{J}(\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})})\big|^{2}+C_{\eta}\big(\|\overline{\partial}_{b}u\|_{k-1}^{2}+\|\overline{\partial}_{b}^{*}u\|_{k-1}^{2}+\|u\|_{k-1}^{2}\big)\;.$
@@
(5.13) $\sum_{j\notin J}\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega_{\eta,\delta}}\big|\alpha_{\eta}(\tilde{L}^{J^{\prime}}_{\chi_{s}T_{\eta}^{k}u_{J}(\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})})\big|^{2}\leq\sum_{j\notin J}\sideset{}{{}^{\prime}}{\sum}_{J}\sideset{}{{}^{\prime}}{\sum}_{|K|=q-1}\int_{b\Omega_{\eta,\delta}}\big|\alpha_{\eta}(\tilde{L}^{K}_{\chi_{s}T_{\eta}^{k}u_{J}(\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})})\big|^{2}\\ \lesssim(1-\eta)\sum_{j\notin J}\sideset{}{{}^{\prime}}{\sum}_{J}\bigg(\|\overline{\partial}_{b}(\chi_{s}T_{\eta}^{k}u_{J}(\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}}))\|^{2}+\|\overline{\partial}_{b}^{*}(\chi_{s}T_{\eta}^{k}u_{J}(\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}}))\|^{2}\bigg)\\ +C_{\eta}\sum_{j\notin J}\sideset{}{{}^{\prime}}{\sum}_{J}\|\chi_{s}T_{\eta}^{k}u(\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})\|_{-1}^{2}$
@@
(5.14) $\sum_{j\notin J}\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega_{\eta,\delta}}\big|\alpha_{\eta}(\tilde{L}^{J^{\prime}}_{\chi_{s}T_{\eta}^{k}u_{J}(\overline{\omega_{j}}\wedge\overline{\omega_{J^{\prime}}})})\big|^{2}\\ \lesssim(1-\eta)\big(\|\overline{\partial}_{b}(\chi_{s}T_{\eta}^{k}u)\|^{2}+\|\overline{\partial}_{b}^{*}(\chi_{s}T_{\eta}^{k}u)\|^{2}+\|\chi_{s}T_{\eta}^{k}u\|^{2}\big)+C_{\eta}\|u\|_{k-1}^{2}\\ \lesssim(1-\eta)\big(\|\chi_{s}\overline{\partial}_{b}(T_{\eta}^{k}u)\|^{2}+\|\chi_{s}\overline{\partial}_{b}^{*}(T_{\eta}^{k}u)\|^{2}+\|\hat{\chi}_{s}T_{\eta}^{k}u\|^{2}\big)+C_{\eta}\|u\|_{k-1}^{2}\;,$
@@
Next, we similarly address the commutator term with $\overline{\partial}_{b}^{*}$ on the right hand side of (5.11). The argument is actually simpler; because $\sum_{j}\chi_{s}(T_{\eta}^{k}u_{jS})L_{j}=\sum_{j}(\chi_{s}T_{\eta}^{k}u)_{jS}L_{j}=\tilde{L}^{S}_{\chi_{s}T_{\eta}^{k}u}$. Formula (3.4) and Proposition 1 give
(5.16) $\|\chi_{s}[\overline{\partial}_{b}^{*},T_{\eta}^{k}]u\|^{2}\\ \lesssim\sideset{}{{}^{\prime}}{\sum}_{S}\big\|\sum_{j}\chi_{s}\alpha_{\eta}(L_{j})T_{\eta}^{k}u_{jS}\big\|^{2}+C_{\eta}\big(\|\overline{\partial}_{b}u\|_{k-1}^{2}+\|\overline{\partial}_{b}^{*}u\|_{k-1}^{2}+\|u\|_{k-1}^{2}\big)\\ =\sideset{}{{}^{\prime}}{\sum}_{S}\|\alpha_{\eta}(\tilde{L}^{S}_{\chi_{s}T_{\eta}^{k}u})\|^{2}+C_{\eta}\big(\|\overline{\partial}_{b}u\|_{k-1}^{2}+\|\overline{\partial}_{b}^{*}u\|_{k-1}^{2}+\|u\|_{k-1}^{2}\big)\\ \lesssim(1-\eta)\big(\|\overline{\partial}_{b}(\chi_{s}T_{\eta}^{k}u)\|^{2}+\|\overline{\partial}_{b}^{*}(\chi_{s}T_{\eta}^{k}u)\|^{2}\big)\;\;\;\;\;\;\;\;\;\;\;\;\;\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+C_{\eta}\big(\|\overline{\partial}_{b}u\|_{k-1}^{2}+\|\overline{\partial}_{b}^{*}u\|_{k-1}^{2}+\|u\|_{k-1}^{2}\big)\\ \;\;\;\;\;\;\;\;\;\lesssim(1-\eta)\big(\|\chi_{s}\overline{\partial}_{b}(T_{\eta}^{k}u)\|^{2}+\|\chi_{s}\overline{\partial}_{b}^{*}(T_{\eta}^{k}u)\|^{2}+\|\hat{\chi}_{s}T_{\eta}^{k}u\|^{2}\big)\\ +C_{\eta}\big(\|\overline{\partial}_{b}u\|_{k-1}^{2}+\|\overline{\partial}_{b}^{*}u\|_{k-1}^{2}+\|u\|_{k-1}^{2}\big)\;.$